[Hallc_running] [New Logentry] RC Weekly Summary

billlee at jlab.org billlee at jlab.org
Mon Apr 15 17:30:03 EDT 2019

Logentry Text:
Dear Collaborators,

This would be the last RC update/summary for the run period.  During the last week of Kaon L/T program, we had very successfully running. Data were collected at two SHMS angle (SHMS=8.2 and SHMS=6.2 ) at Q2=5.5, W=3, low epsilon. Both angle settings give a low rate ~ 5-12 Kaons (through Lambda channel) per hour. We exceeded the projected target in both angle settings, and collected 850+ K Lambda at SHMS=8.2, and 1100+ at SHMS=6.2.

The only down side during last week of running was the loss of 10 hour beam time due to the SHMS dipole led flow valve controller failure during Thursday owl shift, experts and management are looking into procedural solutions the to minimize future loss.

There is no out standing issues, however, small glitches such as: target communication (white) alarm, target fan current register 7500A, infinite raster energy, persisted during the operation. Also, the wifi base station was rebooted hoping to resolve laptop connection issue.

Many thanks


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3679225
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190415/da474206/attachment.html>

More information about the Hallc_running mailing list